# LayerSync: Self-aligning Intermediate Layers

### v1

@@ 4.3 Text-condition Human Motion Generation
Implementation details. We follow the exact setup as MDM  (Tevet et al., 2023) using a transformer with 8 layers. We use HumanML3D dataset (Guo et al., 2022a). We train the model with and without LayerSync for 600K iterations. We align layer 3 with 6. More details are provided in Appendix F
Evaluation metrics. We report FID and R-Percision (top 3)  (Kynkäänniemi et al., 2019) as detailed in Appendix H.
@@
4.4 Representation Learning
To evaluate the effect of LayerSync we analyze the model’s intermediate representations. We compare SiT-XL/2 model trained with LayerSync for 160 epochs against a baseline SiT-XL/2 trained for 1400 epochs as they both have similar FID. This ensures that both models exhibit comparable generative performance, allowing us to isolate the impact of our regularization on the learned representations, independently of the final generation quality.
We consider linear probing for classification on Tiny ImageNet dataset  (Deng et al., 2009), linear probing for segmentation on the PASCAL VOC dataset  (Everingham et al., 2010), and Centered Kernel Alignment (CKA)  (Kornblith et al., 2019) with DINOv2 embeddings  (Oquab et al., 2023) to measure the distance between the model representations. More implementation details are provided in Appendix E.
We conclude that the representational benefits of LayerSync are not merely a byproduct of accelerated convergence. Even when the baseline model is afforded more than 8x larger training budget to match generative performance, its internal representations remain significantly inferior. We therefore hypothetize that LayerSync acts as a powerful structural regularizer that fundamentally alters the model’s optimization trajectory. By imposing an internal semantic constraint, it guides the network to discover a more efficient and globally coherent feature hierarchy, one that remains inaccessible to the unconstrained model.
Furthermore, it is noteworthy that the quality of representations learned with LayerSync approaches that of models trained with powerful external guidance. Our peak classification accuracy, for example, is comparable to the results reported by (Yu et al., 2024). We interpret this as evidence for our initial hypothesis: LayerSync establishes a virtuous cycle that progressively refines the entire feature hierarchy. Notably, the layer of peak performance often shifts to align with the chosen reference layer. For instance, when layer 8 was synced with layer 16, the new performance peak emerged at layer 16. One possible interpretation is that the alignment process redefines the model’s internal structure, effectively positioning the reference layer as the new frontier between feature encoding and decoding. We leave a deeper investigation and further evaluation of LayerSync’s impact on representations as future work.
@@ 5 Ablation Study
Layer Selection. To empirically validate the robustness of our layer selection strategy, we conducted an experiment with randomized layer pairings. For both SiT-XL and SiT-B architectures, we performed 10 independent training each with a different, randomly selected pair of layers following our simple heuristic proposed in Section 3.4. The results in Table 5 demonstrate remarkable consistency. The low standard deviation in the FID (0.8 for SiT-XL) on both architecture confirms that the specific choice of layers is not a very sensitive hyperparameter. This robustness validates our claim that LayerSync is a practical, plug-and-play method that provides significant performance gains without necessitating an expensive search for optimal layer combinations.
Table 5: Performance of LayerSync with randomized layer pairings. Results show the mean FID and standard deviation (in parentheses) over 10 independent runs, confirming the robustness of LayerSync to layer selection.
Method Model Iterations FID $\downarrow$ (STD)
Baseline SiT-B 400k 36.19
Dispersive SiT-B 400k 32.45
LayerSync - Ours SiT-B 400k 31.38 (0.7)
Baseline SiT-XL 400k 17.98
Dispersive SiT-XL 400k 15.59
LayerSync - Ours SiT-XL 400k 12.24 (0.8)
@@
Table 6: Ablation study for $\lambda$. We train SiT B/2 for 400K iterations while aligning block 2 with 8. We observe that our approach is robust for a wide range of $\lambda$. The baseline SiT B/2 has FID 36.19.
@@ 6 Discussion
This similarity between layers raises a natural question: does LayerSync make layers redundant, potentially allowing for model pruning? Our experiments (Appendix B) show that while models trained with LayerSync are more robust to layer removal than their baseline counterparts, performance still degrades significantly. This indicates that despite the improved alignment, each layer retains a unique function essential to the model’s capacity. Consequently, naively pruning a trained model did not prove superior to simply training a smaller architecture from scratch. However, the increase in resilience to layer removal is a finding that suggests that LayerSync may alter the functional contribution of layers in a way that could require further investigation.
@@ Appendix A Video generation
Fine-tuning on SSv2. We use SSv2 dataset  (Goyal et al., 2017) to finetune CogVidX-2B video generation model  (Yang et al., 2024) and Wan2.1 1.3B foundation model  (Wan et al., 2025). Each video has 33 frames. The frames are normalized and processed with Stable Diffusion VAE  (Rombach et al., 2022). Instead of cosine similarity we use the TRD similarity metric proposed in Zhang et al. (2025), as it has been shown that it is more effective than cosine similarity for finetuning. We align layer 4 with layer 24. Each model is fine-tuned for a single epoch (1,100 steps) using 16 GPUs and a global batch size of 160 videos.
TRD loss definition: The loss is based on  (Zhang et al., 2025) and has two terms. The spatial term is a per-frame similarity metric, and the temporal term focuses on the consistency between the frame and is a cross-frame similarity metric. Assuming $\mathbf{y}_{v}$ being the intermediate representation, we first reshape it to $\mathbb{R}^{f\times(hw)\times D}$, with $f$ being the number of frames, $hw$ being the size of the token and $D$ being the representation dimension, then the spatial term is calculated as:
$y^{d,i,j}_{\text{spatial}}=\frac{\mathbf{y}_{v}^{d,i}\cdot\mathbf{y}_{v}^{d,j}}{\|\mathbf{y}_{v}^{d,i}\|\,\|\mathbf{y}_{v}^{d,j}\|},$ (4)
where $i,j\in[1,hw]$ index spatial positions and $d$ being the frame index.
An the temporal term is calculated as:
$y^{d,i,j,e}_{\text{temp}}=\frac{\mathbf{y}_{v}^{d,i}\cdot\mathbf{y}_{v}^{e,j}}{\|\mathbf{y}_{v}^{d,i}\|\,\|\mathbf{y}_{v}^{e,j}\|},\quad\forall e\in[1,f]\setminus\{d\},\;j\in[1,hw].$ (5)
The final TRD loss term is :
$\mathcal{L}_{\text{TRD}}=\underbrace{\frac{1}{f(hw)^{2}}\sum_{d=1}^{f}\sum_{i,j=1}^{hw}\big|h^{d,i,j}_{\text{spatial}}-y^{d,i,j}_{\text{spatial}}\big|}_{\text{Spatial component}}+\underbrace{\frac{1}{f(hw)^{2}(f-1)}\sum_{d=1}^{f}\sum_{e\neq d}\sum_{i,j=1}^{hw}\big|h^{d,i,j,e}_{\text{temp}}-y^{d,i,j,e}_{\text{temp}}\big|}_{\text{Temporal component}},$ (6)
where $h$ can be either an external representation or, in our case the representtaion of a different layer.
Baselines. We compare LayerSync with Dispersive  (Wang & He, 2025). We apply the dispersive loss at 25% depth which is what yielded the best results in original work. We refer to fine-tuning without any extra guidance as vanilla.
Evaluation metrics. We rely on Fréchet Video Distance (FVD; Unterthiner et al. (2018)) for evaluation. We generate 5000 videos of 33 frames for evaluation on finetuned models and 16 frames for SiT-XL.
Results. As shown in Table 7, LayerSync consistently outperforms all baselines across both fine-tuning and from-scratch training setups, achieving the lowest FVD scores in every scenario. When fine-tuning large pre-trained models on SSv2, LayerSync improves FVD by 19.1% over the vanilla baseline for CogVideoX-2B and by 22.8% for Wan2.1, demonstrating its effectiveness in enhancing temporal coherence and sample quality during adaptation. In the from-scratch CLEVRER experiment, LayerSync achieves a 54.7% reduction in FVD compared to the vanilla baseline. This substantial gain highlights LayerSync’s ability to serve as a strong inductive bias, improving learning efficiency and generation quality.
These consistent improvements across model scales and training regimes underscore the generality and robustness of LayerSync as a self-contained regularization strategy.
Table 7: FVD scores ($\downarrow$) for video generation. We observe that LayerSync consistently improves FVD for both finetuning and training from scratch.
CogVideoX-2B (SSv2) Wan2.1 (SSv2) SiT-XL (CLEVERER)
Vanilla 371.88 363.98 265.50
Dispersive 342.10 372.43 165.12
LayerSync (Ours) 300.91 280.78 120.13
@@ Appendix B Dropping Blocks
To investigate the effect of dropping blocks, we train SiT XL/2 and SiT XL/2 with LayerSync (aligning layers 7 and 16) for 120k iterations. We then drop four blocks in between the aligned layers. Quantitative results are presented in Table 8, and qualitative examples are shown in Figure 4, indicating that the model trained with LayerSync is more robust to block drop.
@@
Appendix C LayerSync - Algorithm

### v2

@@ 4.3 Text-conditional Human Motion Generation
Implementation details. We follow the exact setup as MDM  (Tevet et al., 2022) using a transformer with 8 layers. We use HumanML3D dataset (Guo et al., 2022). We train the model with and without LayerSync for 600K iterations. We align layer 3 with 6. More details are provided in Appendix M
Evaluation metrics. We report FID and R-Precision (top 3)  (Kynkäänniemi et al., 2019) as detailed in Appendix O.
@@
4.4 Representation Learning
To evaluate the effect of LayerSync, we analyze the model’s intermediate representations. We compare SiT-XL/2 model trained with LayerSync for 160 epochs against a baseline SiT-XL/2 trained for 1400 epochs as they both have similar FID. This ensures that both models exhibit comparable generative performance, allowing us to isolate the impact of our regularization on the learned representations, independently of the final generation quality.
We consider linear probing for classification on Tiny ImageNet dataset  (Deng et al., 2009), linear probing for segmentation on the PASCAL VOC dataset  (Everingham et al., 2010), and Centered Kernel Alignment (CKA)  (Kornblith et al., 2019) with DINOv2 embeddings  (Oquab et al., 2023) to measure the distance between the model representations. More implementation details are provided in Appendix L.
We conclude that the representational benefits of LayerSync are not merely a byproduct of accelerated convergence. Even when the baseline model is afforded more than 8x larger training budget to match generative performance, its internal representations remain significantly inferior. We therefore hypothesize that LayerSync acts as a powerful structural regularizer that fundamentally alters the model’s optimization trajectory. By imposing an internal semantic constraint, it guides the network to discover a more efficient and globally coherent feature hierarchy, one that remains inaccessible to the unconstrained model.
Furthermore, it is noteworthy that the quality of representations learned with LayerSync approaches that of models trained with powerful external guidance. Our peak classification accuracy, for example, is comparable to the results reported by (Yu et al., 2024). We interpret this as evidence for our initial hypothesis: LayerSync establishes a virtuous cycle that progressively refines the entire feature hierarchy. Notably, the layer of peak performance often shifts to align with the chosen reference layer. A more detailed analysis is provided in Appendix G.
4.5 LayerSync Combined with External Representations
In this experiment, we investigate whether LayerSync can be combined with external representation guidance to further accelerate training. The results summarized in Table 5 show that the two approaches are synergistic; combining LayerSync with REPA  (Yu et al., 2024) yields better performance than REPA alone. This suggests that the internal structural alignment of LayerSync (see Section 3.2) and the external semantic injection of REPA operate as complementary axes of improvement. Thus, LayerSync can be used as a standalone method when external models are unavailable or too costly, or combined with external representations to maximize performance. Additional details provided in Appendix H
Table 5: Quantitative evaluation of SiT-XL/2 trained on ImageNet $256\times 256$ (Deng et al., 2009) comparing REPA and REPA + LayerSync. Both models are trained for 200k iterations and a batch size of 1024. REPA is applied on Layer 10 and LayerSync aligns layer 8 with 16. The results show that LayerSync can be combined with approaches that rely on external representations, such as REPA, to further accelerate training.
Method FID $\downarrow$ sFID $\downarrow$ Inception Score $\uparrow$
REPA 7.88 4.81 126.39
REPA + LayerSync 7.01 4.78 129.85
@@ 5 Ablation Study
Layer Selection. To empirically validate the robustness of our layer selection strategy, we conducted an experiment with randomized layer pairings. For both SiT-XL and SiT-L architectures, detailed in Tables 11 and 12. We observe that regardless of the selected layer, we always have a gain. However, the gain is optimal if following the layer selection strategy discussed in Section 3.4. The results in Table 6 demonstrate remarkable consistency. The low standard deviation in the FID (0.8 for SiT-XL) confirms that the specific choice of layers is not a very sensitive hyperparameter. This robustness validates our claim that LayerSync is a practical, plug-and-play method that provides significant performance gains without necessitating an expensive search for optimal layer combinations.
Table 6: Performance of LayerSync with randomized layer pairings. Results show the mean FID, IS and standard deviation (in parentheses) over independent runs, confirming the robustness of LayerSync to layer selection.
Method Model FID $\downarrow$ (STD) IS $\uparrow$ (STD)
Baseline SiT-XL 17.98 50.41
LayerSync - Ours SiT-XL 12.24 (0.8) 71.47 (2.6)
@@
Table 7: Ablation study for $\lambda$. We train SiT-B/2 for 400K iterations while aligning block 2 with 8. We observe that our approach is robust for a wide range of $\lambda$. The baseline SiT-B/2 has FID 36.19.
@@ 6 Discussion
This similarity between layers raises a natural question: does LayerSync make layers redundant, potentially allowing for model pruning?
Our experiments (Appendix B) show that while models trained with LayerSync are more robust to layer removal than their baseline counterparts, performance still degrades significantly. This indicates that despite the improved alignment, each layer retains a unique function essential to the model’s capacity. Consequently, naively pruning a trained model did not prove superior to simply training a smaller architecture from scratch. Similar findings have been reported for LLMs: intermediate transformer blocks often exhibit high correlation. While removing these blocks has little effect on easier tasks such as question answering (QA), it degrades performance on more challenging tasks (Gromov et al., 2024). Therefore, high inter-block correlation does not necessarily imply that the blocks are redundant.
@@ Appendix A Video generation
Fine-tuning on MixKit dataset. We use MixKit dataset  (Lin et al., 2024) to fine-tune Wan2.1 1.3B foundation model  (Wan et al., 2025). Each video has 81 frames. We align layer 8 with layer 16 with a regularizer weight of 0.02. The model is fine-tuned via LoRA adapters for 1,100 steps using 4 GPUs.
Baselines. We refer to fine-tuning without any extra guidance as vanilla.
Evaluation metrics. We rely on Fréchet Video Distance (FVD; Unterthiner et al. (2018)) for evaluation. We generate 5,000 videos of 16 frames for SiT-XL and 1,000 videos of 81 frames for the fine-tuned Wan2.1 model.
Results. As shown in Table 8, LayerSync consistently outperforms vanilla training across both fine-tuning and training from-scratch setups, achieving the lowest FVD scores in every scenario.
Table 8: FVD scores ($\downarrow$) for video generation. We observe that LayerSync consistently improves FVD for both fine-tuning and training from scratch.
Wan2.1 (MixKit) SiT-XL (CLEVRER)
Vanilla 321.84 265.50
LayerSync (Ours) 304.68 120.13
@@ Appendix B Dropping Blocks
To investigate the effect of dropping blocks, we train SiT XL/2 and SiT XL/2 with LayerSync (aligning layers 7 and 16) for 120k iterations. We then drop four blocks in between the aligned layers. Quantitative results are presented in Table 9, and qualitative examples are shown in Figure 5, indicating that the model trained with LayerSync is more robust to block removal.
@@
Appendix C Alignment Considering the Timestep
We apply alignment on SiTXL/2 for the last 75 %, 50 % and 25 % of timesteps. The model is trained on ImageNet $256\times 256$ for 80 epochs. The results summarized in Table 10 shows that the best performance is achieved when alignment is applied on all the timesteps, which confirms that improving the weak representations is beneficial regardless of the timestep.
Table 10: Applying LayerSync on specific timesteps shows that alignment is beneficial for all the timesteps.
Timestep FID $\downarrow$
25% 18.28
50% 18.77
75% 17.68
100% 16.03
Appendix D Ablation on Block Selection
Ablation study on block selection for SiT-XL/2 and SiT-L/2 summarized in Tables 11 and 12 shows that LayerSync consistently improves the generation quality, but the gain is suboptimal when the distance between the blocks is low or when aligning with the decoder blocks (very last layers). All the models are trained for 100K iterations with 16 GPUs and a batch size of 1024.
Table 11: Ablation study on block selection for SiT-XL/2. The results show that LayerSync consistently improves the generation quality, but the gain is suboptimal when the distance between the blocks is low or when aligning with the decoder blocks (very last layers).
Model IS $\uparrow$ FID $\downarrow$ sFID $\downarrow$
Vanilla SiT-XL/2 50.408 26.534 5.035
LayerSync $6\leftarrow 18$ 75.221 15.386 4.672
LayerSync $7\leftarrow 17$ 75.727 15.433 4.694
LayerSync $7\leftarrow 15$ 74.987 15.740 4.611
LayerSync $9\leftarrow 17$ 73.791 16.058 4.695
LayerSync $8\leftarrow 18$ 73.549 16.078 4.589
LayerSync $8\leftarrow 17$ 72.821 16.276 4.638
LayerSync $8\leftarrow 20$ 71.477 16.568 4.662
LayerSync $9\leftarrow 20$ 71.016 16.661 4.695
LayerSync $8\leftarrow 19$ 71.094 16.680 4.658
LayerSync $6\leftarrow 14$ 72.750 16.697 4.695
LayerSync $10\leftarrow 18$ 70.209 16.999 4.638
LayerSync $9\leftarrow 19$ 69.831 17.091 4.701
LayerSync $11\leftarrow 21$ 68.952 17.096 4.673
LayerSync $7\leftarrow 21$ 69.438 17.126 4.685
LayerSync $11\leftarrow 19$ 68.464 17.615 4.628
LayerSync $9\leftarrow 22$ 68.127 17.823 4.717
LayerSync $8\leftarrow 21$ 67.633 18.032 4.705
LayerSync $6\leftarrow 8$ 57.956 22.630 4.903
LayerSync $6\leftarrow 9$ 56.552 23.307 4.892
LayerSync $6\leftarrow 7$ 55.717 23.924 4.954
LayerSync $11\leftarrow 23$ 61.851 19.966 4.769
LayerSync $12\leftarrow 23$ 59.692 20.711 4.737
LayerSync $15\leftarrow 23$ 55.235 23.030 4.783
Table 12: Ablation study on block selection for SiT-L/2. The results show that LayerSync consistently improves the generation quality, but the gain is suboptimal when the distance between the blocks is low or when aligning with the decoder blocks (very last layers).
Model IS FID sFID
LayerSync $6\leftarrow 15$ 64.638 19.165 4.827
LayerSync $8\leftarrow 16$ 62.794 19.566 4.856
LayerSync $7\leftarrow 15$ 62.854 19.663 4.841
LayerSync $6\leftarrow 14$ 62.616 19.945 4.791
LayerSync $5\leftarrow 13$ 62.158 20.048 4.778
LayerSync $7\leftarrow 16$ 61.4 20.055 4.865
LayerSync $7\leftarrow 18$ 60.533 20.324 4.820
LayerSync $6\leftarrow 16$ 61.304 20.444 4.794
LayerSync $5\leftarrow 12$ 60.807 20.492 4.775
LayerSync $5\leftarrow 11$ 61.104 20.696 4.821
LayerSync $5\leftarrow 15$ 60.362 20.734 4.824
LayerSync $5\leftarrow 17$ 61.178 20.771 4.878
LayerSync $5\leftarrow 8$ 50.473 26.511 5.084
LayerSync $5\leftarrow 7$ 49.948 27.004 5.171
LayerSync $5\leftarrow 6$ 47.19 28.741 5.21
LayerSync $11\leftarrow 20$ 43.253 30.57 5.172
LayerSync $17\leftarrow 20$ 43.563 31.056 5.308
LayerSync $16\leftarrow 20$ 43.242 31.24 5.316
LayerSync $13\leftarrow 20$ 42.535 31.407 5.182
LayerSync $14\leftarrow 20$ 41.626 32.35 5.282
LayerSync $15\leftarrow 20$ 41.381 32.622 5.403
Appendix E LayerSync vs. Increasing Learning Rate
To show that LayerSync impact is not simply due to an increase in the gradient magnitude, we designed two sets of experiments below. All the models are trained with 16 GPUs, batch size 1024 for 100k iterations (80 epochs). We then report the gradient norms and the FIDs of different configurations. We consistently observed that the gradient norm with LayerSync is actually smaller than that of the baseline, indicating that our method is not simply equivalent to using a larger learning rate.
(1) Global learning rate increase.
Starting from the default learning rate of $1\times 10^{-4}$, we trained models with higher learning rates $2\times 10^{-4}$ and $5\times 10^{-4}$. For learning rates above $5\times 10^{-4}$, the model diverges. While increasing the learning rate can partially accelerate training, the resulting FID scores remain worse than those obtained with LayerSync with the default learning rate $1\times 10^{-4}$ as shown in Table 13. The visualization of gradient norm is provided in Figure 7(a).
Table 13: Effect of Global Learning Rate. Simply increasing the global learning rate improves FID slightly but does not match the performance gains of LayerSync.
Method lr. FID
SiT-XL/2 $1\times 10^{-4}$ 26.53
SiT-XL/2 $2\times 10^{-4}$ 24.95
LayerSync $1\times 10^{-4}$ 16.03
(2) Higher learning rate on early blocks only.
We then increased the learning rate only for the first 8 blocks and compared this to a model trained with LayerSync aligning layers 8–16 at a global learning rate of $1\times 10^{-4}$. Again, for learning rates above $1\times 10^{-3}$ training diverges. As summarized in Table 14, for $2\times 10^{-4}$, $5\times 10^{-4}$, and $1\times 10^{-3}$, the FID improvements do not match those obtained with LayerSync. The visualization of the gradient norms is provided in Figure 7(b).
Table 14: Effect of Early-Layer Learning Rate. Increasing the learning rate specifically on early layers (first 8 blocks) is beneficial but still underperforms compared to LayerSync.
Method lr. Early Layers General lr. FID
SiT-XL/2 $\times 10^{-4} 1×1$$\times 10^{-4}$ 26.53
SiT-XL/2 $2\times 10^{-4}$ $1\times 10^{-4}$ 19.24
SiT-XL/2 $5\times 10^{-4}$ $1\times 10^{-4}$ 24.63
LayerSync $\times 10^{-4} 1×1$$\times 10^{-4}$ 16.03
(a) Gradient norm for different learning rates.
(b) Gradient norm when changing the learning rate of early blocks.
Figure 7: Gradient norm visualization revealed that LayerSync impact is not similar to simply increasing the learning rate as the model trained with LayerSync has lower gradient norm than the baseline.
Appendix F LayerSync vs Self-Representation Alignment using EMA
We compare our LayerSync approach with the concurrent work SRA (Jiang et al., 2025) in terms of training time and computational overhead. We computed the metrics for SiT-XL/2 using 4 GH200 GPUs and a batch size of 32 per GPU. Results are reported in Table 15. We show that LayerSync requires $25.5\%$ fewer Flops, is $40.5\%$ faster in real-time, and reaches an FID $5\%$ higher.
Table 15: Comparison between LayerSync and SRA. LayerSync results in lower FID while being less computationally expensive.
Method FID $\downarrow$ Wall-clock time/step $\downarrow$ GFlops $\downarrow$
SiT-XL/2+SRA 1.58 0.617 30762
SiT-XL/2+LayerSync 1.49 0.367 22910
Appendix G The Virtuous Cycle and the Evolution of Representations
To provide empirical support for the ”virtuous cycle” and analyze the evolution of the feature hierarchy, we conducted two additional studies evaluating the segmentation performance (mIoU) of internal representations throughout the network on PASCAL VOC (Everingham et al., 2010).
First, we tracked the evolution of representations throughout the training process (from 100k to 600k steps). As shown in Figure 8(a), we observe that while the relative structure of feature quality across layers remains stable, the performance monotonically improves across the entire hierarchy.
Our second study evaluates models trained with different alignment targets at the same training stage (100k steps) (see Figure 8(b)). This comparison provides the most compelling evidence for our hypothesis. When we synchronize an early block (e.g., block 6) with a deeper target (block 18 vs. block 10), we observe two critical effects:
  • Global Improvement: The model guided by the deeper layer achieves superior downstream performance and lower FID (15.39 vs. 23.90), indicating that guidance from deeper layers correlates with better overall representations.
  • Accelerated Maturation: Notably, using a deeper target shifts the peak performance of the network to earlier layers. By effectively ”pulling” semantic richness from the deep target to the earlier block, the early layers appear to acquire higher-level features sooner in the depth hierarchy.
These observations are consistent with the hypothesized virtuous cycle: guiding an early block with a stronger target improves its representation, which in turn provides higher-quality input to subsequent layers. This likely facilitates the learning of stronger deep representations, which then serve as even better guides, progressively refining the entire hierarchy.
(a) Evolution of representations throughout the training process assessed by semantic segmentation on PASCAL VOC dataset.
(b) Evaluation of representations for different layers alignments assessed by semantic segmentation on PASCAL VOC dataset.
Figure 8: Evaluation of the representations assessed by semantic segmentation on PASCAL VOC dataset. (a) Evolution throughout the training process. (b) Evaluation of representation for different layer alignment.
Appendix H Optimal Placement of External Guidance when Combined with LayerSync
To maximize the synergy between internal and external alignment, we investigated the optimal depth for applying REPA. The results are summarized in Table 16 and show that applying REPA before the synchronization range leads to no significant synergies. The most effective strategy integrates the external signal between the aligned layers. For instance, using LayerSync to align layers 8 and 16 while applying REPA at layer 10 yields the best performance. All the models are trained for 50k iterations with 16 GPUs and a batch size of 1024. To show that the trend continues, we continue training the models to 200k iterations and report the results in the main paper.
Table 16: Qualitative comparison between different combinations of REPA and LayerSync. The results show that combining LayerSync with REPA can further accelerate the training, and the best place to apply REPA is between the syncing layers.
Method REPA Layer LayerSync Layer FID $\downarrow$
SiTXL/2 59.45
SiTXL/2 + LayerSync 8–16 46.26
SiTXL/2 + REPA 7 46.06
SiTXL/2 + REPA + LayerSync 7 8–16 43.55
SiTXL/2 + REPA + LayerSync 10 8–16 29.68
Appendix I Evolution of Block Structure
A comparison between the block structure of the SiT-XL/2 and SiT-XL/2 + LayerSync is provided in Figure 9, showing that LayerSync imposes the structural equilibrium early in training.
(a) Correlation between the blocks of SiT-XL/2 after 120k iterations.
(b) Correlation between the blocks of SiT-XL/2 with LayerSync after 120k iterations.
Figure 9: Evolution of the Block Structure. (a) Early in training (120k iterations), the vanilla model lacks this structure. (b) LayerSync imposes such structural equilibrium early in training (120k iterations).
Appendix J LayerSync - Algorithm